OpenStudy (anonymous):

To convert 132 milligrams to grams, you can set up the proportion x/132mg =1,000mg/1g true or false?

jimthompson5910 (jim_thompson5910):

1 g = 1000 mg (1 g)/1000 = (1000 mg)/1000 0.001 g = 1 mg 1 mg = 0.001 ------------------------------------------------------- 1 mg = 0.001 g 132*(1 mg) = 132*(0.001) g 132 mg = 0.132 g So there are 0.132 grams in 132 milligrams If the given equation was true, then x should be x = 0.132 Let's see if that's the case x/132 = 1000/1 x/132 = 1000 x = 1000*132 x =132000 So the answer is way off. Therefore, the proportion is not the correct proportion to convert 132 milligrams to grams.

OpenStudy (anonymous):

True!